OpenStudy (anonymous):

Solve the system of linear equations below. 6x + 3y = 33 4x + y = 15

OpenStudy (anonymous):

substitute the value of y in the second equation from the first

OpenStudy (anonymous):

6x + 3y = 33 4x + y = 15 | *-3 Can you do it now?

OpenStudy (anonymous):

6x + 3y = 33.............(1) 4x + y = 15...............(2) multiply (2) by 3 3*(4x + y = 15) 12x + 3y = 45..........(3) 6x + 3y = 33 subtract 6x = 12 x=2 substitute x = 2 in an equation and solve for Y y =7